OpenStudy (anonymous):

Steps to solving: 8x - 2y + z= -6 -x + 3y - 2z= -15 3x - y + 4z= 13

OpenStudy (anonymous):

Multiply the second equation with 8 then add the first two equation x will get cancelled let me know what you get after addition.

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

i got 22y-15z=-126

OpenStudy (anonymous):

yes this is correct.

OpenStudy (anonymous):

ok now multiply again the second equation with 3 and add this with the equation 3. multiply the second equation this one with 3 -x + 3y - 2z= -15 then add it with 3x - y + 4z= 13 let me know what you get after addition

OpenStudy (anonymous):

i got 8y-2z=-32

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now you have two equations with two unknowns, y and z 22y-15z=-126 8y-2z=-32 can you find y and z from these equations ??

OpenStudy (anonymous):

using the elimination process i multiplied the new #1 equation by by 8 and the 2nd new equation by 2 and got : -176y+120z=1008 (+)176y-44z=-64 ______________________ and got z=12.4

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

if so, then just plug-in z to get x and y :)

OpenStudy (anonymous):

i did and got x= -5.9 & y= -14.4 but when i plug it in into one of the original equation the values don't work.

OpenStudy (anonymous):

Actually wendy, you should try solving again. I got 4 for z.

OpenStudy (anonymous):

wait yes it did work i check it .

OpenStudy (anonymous):

So tell me what you have for x and y if z=4.

OpenStudy (anonymous):

oh but how come then it works because i also used a linear equation solver in the internet and it also had z=4 so that's why im confused and i got -5.9 for x and -14.4 for y

OpenStudy (anonymous):

It's because when you were eliminating, you multiplied 2 instead of 22. =P check your work.

OpenStudy (anonymous):

using the elimination process i multiplied the new #1 equation by by 8 and the 2nd new equation by 2 and got : -176y+120z=1008 (+)176y-44z=-64 ______________________ and got z=12.4 See that's what you did. You were supposed to get 704 as a constant. Not -64

OpenStudy (anonymous):

sorry guys i was doing my little work:P so where we are ?

OpenStudy (anonymous):

@wendy67 is trying to figure out why she isn't getting z=4....

OpenStudy (anonymous):

22y-15z=-126 8y-2z=-32 did you solve it ?

OpenStudy (anonymous):

hmm.......

OpenStudy (anonymous):

ok let me do this.

OpenStudy (anonymous):

she eliminated wrong

OpenStudy (anonymous):

in the system you gave her @sami-21

OpenStudy (anonymous):

when i did this 22y-15z=-126 8y-2z=-32 i multiplied the eq#1 by 8 and eq#2 by 22 and when i mulitplied the sencond equation by 22 the constant is 704 but i got -704

OpenStudy (anonymous):

-176y+120z=1008 (+)176y-44z=-704nd ______________________ so z=4

OpenStudy (anonymous):

solve equation two 8y-2z=-32 for z -2z=-32-8y z=16+4y and put in 1 22y-15z=-126 tell me what you get for y.

OpenStudy (anonymous):

@wendy67

OpenStudy (anonymous):

so instead of solving for z solve for y?

OpenStudy (anonymous):

ok let me do this solve equation two 8y-2z=-32 for z -2z=-32-8y z=16+4y put this value of z in 22y-15z=-126 22y-15(16+4y)=-126 22y-240-60y=-126 -44y=-126+240 -38y=114 y=-114/38 y=-3 now put this in any of the above two equations i am putting in 8y-2z=-32 8(-3)-2z=-32 -24-2z=-32 -2z=-32+24 -2z=-8 z=4 now use these value of y and z in any of the , original 3 equations given in the question to get x.

OpenStudy (anonymous):

i hope it is clear now :)

OpenStudy (anonymous):

ok i figuered out what i did wrong so i took z=4 and pluged it into the 2nd original equation -x=3y-2(4)=-15 -and got x=7

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

nopes ! look at what i got for y and z then using these value find the value of x

OpenStudy (anonymous):

x=-2?

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

ok to get myself clearer when i solve for z which 4 then i plug it in into and original equation and solve for y and whatever i get for y plug it back in into and original equatin and solve for x?

OpenStudy (anonymous):

*and=an* srry

OpenStudy (anonymous):

nevermind i re-did it following through everything we all discussed and got it cleared thanks @sami-21 and @xThatOneKidx :)

OpenStudy (anonymous):

you are welcome:)